lilelac:

Isabel has a previous balance of $872 on a credit card with a 17.8% APR compounded monthly. If she made a payment of $50 this month, what is the new balance on her credit card?

NOCH:

First calculate the interest Interest=previous balance×interest rate You divide the interest rate by 12 months to find the interest For only one month Also convert the interest rate into decimal form 17.8/100= 0.178 Interest=822×(0.178÷12months) Interest=12.19 New balance Previous balance+interest-payment New balance=872+12.93−50=834.93 your answer is: 834.93